# CursorCore: Assist Programming through Aligning Anything

### v1

@@ 1 Introduction
Since the rise of large language models (LLMs), AI-assisted programming technology has developed rapidly, with many powerful LLMs being applied in this field Zan et al. (2022); Liang et al. (2024); Yang et al. (2024). The technology mainly takes two forms. One form involves completing a specified code snippet at the end or inserting corresponding code at a designated position, typically accomplished by foundation models that support relevant input formats Chen et al. (2021); Bavarian et al. (2022). The other form involves generating or editing code snippets based on natural language instructions or reflections through interaction with the environment, usually carried out by instruction models that have been further aligned Shinn et al. (2023); Cassano et al. (2023b); Muennighoff et al. (2024); Paul-Gauthier (2024). Figure 1 shows simple examples of these forms.
@@ 2 Assistant-Conversation: New Conversation Framework for Programming Assistants
In this section, we introduce a new conversational framework, Assistant-Conversation, aimed at simplifying the programming process. The framework leverages all available information during programming to streamline work for programmers.
@@ 2.1 Framework Formulation
Figure 2: Examples of Assistant-Conversation from our training data.
@@ 2.3 Specifications and Implementation
To represent a piece of code like C, we can either use it directly or wrap it in a markdown code block. However, representing code changes, such as H or changes in A, is more complex. We can either use the whole code, patches that alter the code, or records of both the modification locations and the specific changes. Some methods work well but experience issues when handling longer texts, such as outputting the entire modified code, which can be slow. Other methods output minimal content, like providing only the modification locations and changes. These are faster but still not optimal in terms of performance. We represent code changes in the experiments of the main body using the whole code format, and we investigate different ways to represent these modifications, as detailed in Appendix B. Additionally, we explore methods for compressing historical code changes in Appendix F.
In some cases, programmers assign assistants to focus on specific areas of code. They might use the cursor to mark a general location or directly select a range of code, as shown in Figure 2. We handle this by treating them as special tokens (see Appendix D for further details).
@@ 3.1 Benchmark overview
Table 1: APEval Statistics.
@@ C Statistics Mean / Max
As discussed in Section 2.1, programming assistance can involve different types of information, with H and U being optional. Thus, there are four possible combinations of information: H, C, U; H, C; C, U; and only C. HumanEval Chen et al. (2021) is a well-known benchmark for evaluating code completion. It has been extended to assess other tasks such as code insertion Bavarian et al. (2022), instruction-based tasks CodeParrot (2023); Muennighoff et al. (2024), and multilingual generation Zheng et al. (2023b); Cassano et al. (2023a). We refer to these works and further extend it to comprehensively evaluate the model’s ability to assist programming. We randomly categorize each task into one of the four types, then manually implement the functions and simulate the potential instructions that programmers might give to an LLM during the process, collecting all interactions. We invite programmers with varying levels of experience to annotate the data. After processing, we get the new benchmark, Assist Programming Eval (APEval). Detailed statistics are shown in Table 1.
@@ 3.2 Evaluation Process and Metrics
In all tasks, we use the classic Pass@1 metric to execute the generated code, which is the simplest version of the Pass@k metric Chen et al. (2021). Since APEval is an extension of HumanEval, we use the test set created by EvalPlus Liu et al. (2023). We report the results from both the basic and extra tests. We provide the model with relevant information during the programming process, and the model immediately returns the modified code. Some methods may improve performance by increasing the number of output tokens to model the thinking process; we discuss this further in Appendix E.
@@ 4.1 Data Sources
For each code snippet, we use LLMs to generate the corresponding coding history. Since human coding approaches vary widely, we utilize several LLMs, each guided by three distinct prompts, representing novice, intermediate, and expert programmers. The LLMs then return their version of the coding process. Prompts used are shown in Appendix K.
@@ 4.2 Data Processing
In addition to the above, we model selected code regions, cursor positions, and make LLMs create chat-style interactions with users. Further details are provided in Appendix C.
@@ 5.1 Base models
We fine-tune existing base LLMs to assist with programming tasks. Over the past few years, many open-source foundation models have been trained on large code corpora sourced from GitHub and other platforms, demonstrating strong performance in coding. We choose Deepseek-Coder Guo et al. (2024a), Yi-Coder AI et al. (2024) and Qwen2.5-Coder Hui et al. (2024) series, which we refer to as CursorCore-DS, CursorCore-Yi and CursorCore-QW2.5 series after training. Deepseek-Coder has achieved state-of-the-art performance on numerous coding-related benchmarks over the past year, gaining wide recognition. Yi-Coder and Qwen2.5-Coder are the most recently released models at the start of our experiments and show the best performance on many benchmarks for code now. These models are widely supported by the community, offering a good balance between size and performance, making them suitable for efficient experimentation. For ablation experiments, we use the smallest version, Deepseek-Coder-1.3B, to accelerate the process. We use a chat template adapted from ChatML OpenAI (2023) to model Assistant-Conversation during training, as detailed in Appendix I.
@@ 5.2 Training data
We use Programming-Instruct to collect data. For AIprogrammer, we gather code snippets from datasets such as the stack Kocetkov et al. (2023) and oss-instruct Wei et al. (2023b), then prompt LLMs to generate the programming process. For Git commit data, we collect relevant information from editpackft Cassano et al. (2023b) (a filtered version of commitpackft Muennighoff et al. (2024)) and further refine it through post-processing and filtering. Regarding online submission data, we source the programming process from the Codenet dataset Puri et al. (2021). First, we group all submissions by user for each problem, then exclude invalid groups without correct submissions to obtain complete programming processes. These are then fed into the processing pipeline to generate the final training data. In total, we accumulate 219K samples, with detailed statistics and distributions shown in Tables 2, 3, 6, 6, 8 and 8. To ensure compatibility with previous chatbot-style interactions and further improve model performance, we also incorporate the evol-instruct dataset ISE-UIUC (2023) collected using the GPT series Ouyang et al. (2022), which has been widely recognized for its high quality during training. Following StarCoder’s data processing approach Li et al. (2023), we decontaminate our training data.
@@ 6.1 Experimental setup
We conduct the data selection ablation and primary evaluation on our APEval benchmark, and provide results on well-known benchmarks such as Python program synthesis, automated program repair, and instructional code editing, which are detailed in Appendix G. We choose prominent open-source and closed-source LLMs as our baselines. For all benchmarks, we use greedy decoding to generate evaluation results. CursorCore natively supports various inputs in APEval, whereas base and instruction LLMs require additional prompts for effective evaluation. We design few-shot prompts separately for base and instruction models, as detailed in Appendix J.
@@ 6.3 Evaluation results on APEval
In Table 4, we present the results of evaluating CursorCore series models and other LLMs on APEval. It includes both the average results and the results across four different types of information within the benchmark, each item in the table is the score resulting from running the base tests and extra tests. We also report the evaluation results of other well-known models on APEval, which can be found in Appendix H.
@@
In 1B+ parameter models, general LLMs significantly underperform code LLMs. Even the best-performing general model scores over 10% lower on APEval compared to the best-performing code model, despite having more parameters. For models with 6B+ parameters, while general LLMs still lag behind code LLMs, the performance gap narrows considerably, with general LLMs even surpassing in certain cases involving specific information types. When it comes to 10B+ models, the performance difference between general and code LLMs becomes negligible. We think that smaller models, due to their limited parameter capacity, tend to focus on a single domain, such as programming assistance, while larger models can encompass multiple domains without compromising generalizability.
@@ A.1 AI-Assisted Programming
AI-assisted programming has a long history, encompassing various tasks such as clone detection Lu et al. (2021), code summarization Sun et al. (2024), program synthesis Chen et al. (2021); Austin et al. (2021), automatic program repair Gulwani et al. (2016), code editing Wei et al. (2023a), and code optimization Shypula et al. (2024). In the past, these tasks were typically addressed by custom-built models, which were difficult to scale across different tasks. With the rise of LLMs, AI-assisted programming increasingly leverages LLMs to handle multiple types of tasks simultaneously. Numerous high-quality open-source and closed-source products, such as Continue Continue-Dev (2024), Aider Paul-Gauthier (2024), Copilot Github-Copilot (2022) and Cursor Cursor-AI (2023), are based on this approach.
@@ Appendix B Code modification representation
Appendix C Additional details about Programming-Instruct
@@
In addition to generating H and U, as discussed in Section 4.2, we also simulate the programmer’s specification of the target area and model interactions in a chat-style format. The target modification area is created using a random algorithm, as described in Appendix D, while the chat-style interaction is generated using LLMs which is similar to the generation of instructions. Prompts used for it are provided in Appendix K.
Appendix D Target area representation
@@
Appendix E Discussion about thought process
@@
Appendix F Conversation retrieval for Assistant-Conversation
@@
Appendix G Evaluation results of other benchmarks
@@
Appendix H Additional evaluation results on APEval
@@
Appendix I Chat template
@@
Appendix J Prompts for evaluation
@@
Appendix K Prompts for data collection
We design specific system prompts and few-shot examples to collect high-quality training data, as we find that many examples are very difficult to complete with current LLMs, and only a few of them can be successfully completed using rough prompts. For AIprogrammer, we utilize LLMs to simulate programmers at three different skill levels, with each level using a distinct set of prompts as shown in Tables 7, 8 and 9. Additionally, prompts used for evaluating whether the outputs align with user intent, generating user instructions, and facilitating chat interactions between models and users are outlined in Tables 12, 10 and 11. Partial few-shot examples are shown in Figures 18, 19, 20, 21, 22 and 23.
Appendix L Limitations and future work

### v2

@@ 1 Introduction
Since the rise of large language models (LLMs), AI-assisted programming technology has developed rapidly, with many powerful LLMs being applied in this field Zan et al. (2022); Liang et al. (2024); Yang et al. (2024). The technology mainly takes two forms. One form involves completing a specified code snippet at the end or inserting corresponding code at a designated position, typically accomplished by foundation models Chen et al. (2021); Bavarian et al. (2022) that support relevant input formats. The other form involves generating or editing code snippets based on natural language instructions or reflections through interaction with the environment, usually carried out by instruction models that have been further aligned Shinn et al. (2023); Cassano et al. (2023b); Muennighoff et al. (2024); Paul-Gauthier (2024). Figure 1 shows simple examples of these forms.
@@ 2 Assistant-Conversation: New Conversation Framework for Programming Assistants
In this section, we introduce a new conversational framework, Assistant-Conversation, aimed at simplifying the programming process. The framework leverages all available information during programming to streamline work for programmers. By precisely defining various types of information and their formats, Assistant-Conversation directly aligns with the input and output requirements of applications such as automated editing and inline chat. This framework facilitates model alignment, enabling fast and accurate generation and parsing.
@@ 2.1 Framework Formulation
Figure 2: Examples of Assistant-Conversation from our training data. The top example demonstrates predicting the corresponding edits and explanations based on historical edits and the current code. The bottom example demonstrates predictions based on the current code and user instructions.
@@ 2.3 Specifications and Implementation
To represent a piece of code like C, we can either use it directly or wrap it in a markdown code block. However, representing code changes, such as H or changes in A, is more complex. We can either use the whole code, patches that alter the code, or records of both the modification locations and the specific changes. Some methods work well but experience issues when handling longer texts, such as outputting the entire modified code, which can be slow. Other methods output minimal content, like providing only the modification locations and changes. These are faster but still not optimal in terms of performance. We represent code changes in the experiments of the main body using the whole code format, and we investigate different ways to represent these modifications, as detailed in Appendix B. Additionally, we explore methods for compressing historical code changes in Appendix G.
In some cases, programmers assign assistants to focus on specific areas of code. They might use the cursor to mark a general location or directly select a range of code, as shown in Figure 2. We handle this by treating them as special tokens (see Appendix E for further details).
@@ 3.1 Benchmark overview
Table 1: APEval Statistics. Present statistical information about H, C, and U in our benchmark.
@@ C Statistics Mean / Max
As discussed in Section 2.1, programming assistance can involve different types of information, with H and U being optional. Thus, there are four possible combinations of information: H, C, U; H, C; C, U; and only C. HumanEval Chen et al. (2021) is a well-known benchmark for evaluating code completion. It has been extended to assess other tasks such as code insertion Bavarian et al. (2022), instruction-based tasks CodeParrot (2023); Muennighoff et al. (2024), and multilingual generation Zheng et al. (2023b); Cassano et al. (2023a). We refer to these works and further extend it to comprehensively evaluate the model’s ability to assist programming. We randomly categorize each task into one of the four types, then manually implement the functions and simulate the potential instructions that programmers might give to an LLM during the process, collecting all interactions. We invite programmers with varying levels of experience to annotate the data. After processing, we get the new benchmark, Assist Programming Eval (APEval). Detailed statistics are shown in Table 1. Specific details regarding the collection process and examples of our benchmark can be found in Appendix C.
@@ 3.2 Evaluation Process and Metrics
In all tasks, we use the classic Pass@1 metric to execute the generated code, which is the simplest version of the Pass@k metric Chen et al. (2021). Since APEval is an extension of HumanEval, we use the test set created by EvalPlus Liu et al. (2023). We report the results from both the basic and extra tests. We provide the model with relevant information during the programming process, and the model immediately returns the modified code. Some methods may improve performance by increasing the number of output tokens to model the thinking process; we discuss this further in Appendix F.
@@ 4.1 Data Sources
For each code snippet, we use LLMs to generate the corresponding coding history. Since human coding approaches vary widely, we utilize several LLMs, each guided by three distinct prompts, representing novice, intermediate, and expert programmers. The LLMs then return their version of the coding process. Prompts used are shown in Appendix L.
@@ 4.2 Data Processing
In addition to the above, we model selected code regions, cursor positions, and make LLMs create chat-style interactions with users. Further details are provided in Appendix D.
@@ 5.1 Base models
We fine-tune existing base LLMs to assist with programming tasks. Over the past few years, many open-source foundation models have been trained on large code corpora sourced from GitHub and other platforms, demonstrating strong performance in coding. We choose the base versions of Deepseek-Coder Guo et al. (2024a), Yi-Coder AI et al. (2024) and Qwen2.5-Coder Hui et al. (2024) series, as fine-tuning is generally more effective when applied to base models rather than instruction models. After training, we refer to them as CursorCore-DS, CursorCore-Yi and CursorCore-QW2.5 series. Deepseek-Coder has achieved state-of-the-art performance on numerous coding-related benchmarks over the past year, gaining wide recognition. Yi-Coder and Qwen2.5-Coder are the most recently released models at the start of our experiments and show the best performance on many benchmarks for code now. These models are widely supported by the community, offering a good balance between size and performance, making them suitable for efficient experimentation. For ablation experiments, we use the smallest version, Deepseek-Coder-1.3B, to accelerate the process. We use a chat template adapted from ChatML OpenAI (2023) to model Assistant-Conversation during training, as detailed in Appendix J.
@@ 5.2 Training data
We use Programming-Instruct to collect data. For AIprogrammer, we gather code snippets from datasets such as the stack Kocetkov et al. (2023) and oss-instruct Wei et al. (2023b), then prompt LLMs to generate the programming process. For Git commit data, we collect relevant information from editpackft Cassano et al. (2023b) (a filtered version of commitpackft Muennighoff et al. (2024)) and further refine it through post-processing and filtering. Regarding online submission data, we source the programming process from the Codenet dataset Puri et al. (2021). First, we group all submissions by user for each problem, then exclude invalid groups without correct submissions to obtain complete programming processes. These are then fed into the processing pipeline to generate the final training data. In total, we accumulate 219K samples, with detailed statistics and distributions shown in Tables 2, 3, 6, 6, 8 and 8. AIprogrammer data has the shortest average length, while Online Submit data has the longest. To ensure compatibility with previous chatbot-style interactions and further improve model performance, we also incorporate the evol-instruct dataset ISE-UIUC (2023) collected using the GPT series Ouyang et al. (2022), which has been widely recognized for its high quality during training. Following StarCoder’s data processing approach Li et al. (2023), we decontaminate our training data.
@@ 6.1 Experimental setup
We conduct the data selection ablation and primary evaluation on our APEval benchmark, and provide results on well-known benchmarks such as Python program synthesis, automated program repair, and instructional code editing, which are detailed in Appendix H. We choose prominent open-source and closed-source LLMs as our baselines. For all benchmarks, we use greedy decoding to generate evaluation results. CursorCore natively supports various inputs in APEval, whereas base and instruction LLMs require additional prompts for effective evaluation. We design few-shot prompts separately for base and instruction models, as detailed in Appendix K.
@@ 6.3 Evaluation results on APEval
In Table 4, we present the results of evaluating CursorCore series models and other LLMs on APEval. It includes both the average results and the results across four different types of information within the benchmark, each item in the table is the score resulting from running the base tests and extra tests. We also report the evaluation results of other well-known models, which can be found in Appendix I.
@@
In 1B+ parameter models, general LLMs significantly underperform code LLMs. Even the best-performing general model scores over 10% lower compared to the best-performing code model, despite having more parameters. For models with 6B+ parameters, while general LLMs still lag behind code LLMs, the performance gap narrows considerably, with general LLMs even surpassing in certain cases involving specific information types. When it comes to 10B+ models, the performance difference between general and code LLMs becomes negligible. We think that smaller models, due to their limited parameter capacity, tend to focus on a single domain, such as programming assistance, while larger models can encompass multiple domains without compromising generalizability.
@@ A.1 AI-Assisted Programming
AI-assisted programming has a long history, encompassing various tasks such as clone detection Lu et al. (2021), code summarization Sun et al. (2024), program synthesis Chen et al. (2021); Austin et al. (2021), automatic program repair Gulwani et al. (2016), code editing Wei et al. (2023a), and code optimization Shypula et al. (2024). These tasks attempt to incorporate a wide range of information into their processes, such as historical edits Gupta et al. (2023); Zhang et al. (2022) and user instructions Cassano et al. (2023b). In the past, however, they were typically addressed by custom-built models, which were difficult to scale across different tasks and types of information. With the rise of LLMs, AI-assisted programming increasingly leverages LLMs to handle multiple types of tasks simultaneously. Numerous high-quality open-source and closed-source products, such as Continue Continue-Dev (2024), Aider Paul-Gauthier (2024), Copilot Github-Copilot (2022) and Cursor Cursor-AI (2023), are based on this approach.
@@ Appendix B Code modification representation
Appendix C Details regarding the collection process of APEval
We inform the annotators about the function’s entry point and its purpose, and allow them to send instructions to the AI programming assistant at appropriate moments. We then use screen recording tools to capture the annotators’ process of wrtining this function. Afterward, we manually analyze the recordings to construct our benchmark. The historical information, current code, and user instructions are all provided by annotators based on the specified function functionality, to cover various code editing scenarios.
During the process of creating the benchmark, in order to better evaluate the model’s ability to utilize historical edits and integrate this information with user instructions, we collected samples for the (H, C) and (H, C, U) types that required the use of relevant historical information to accurately infer user intent. If a sample contained only a single type of information (such as only C or only U), it might be impossible to provide an adequate answer due to a lack of sufficient information.
In our benchmark collection process, we initially annotated one programming process for each task. For some tasks, the annotators consulted the programming assistant; for others, they did not. Similarly, some tasks involved complex editing histories, while others did not. Upon reviewing the data, we found that for certain tasks, it was nearly impossible to collect realistic programming processes containing specific types of information. For example, Some tasks are straightforward and can be completed with just a few lines of code. Programmers who have undergone basic training can write these solutions quickly without needing to consult an assistant or repeatedly revise their code. Conversely, some tasks may involve calling specific libraries or algorithms that most annotators are unfamiliar with, leading them to rely on the programming assistant. It would be unrealistic and counterproductive to instruct annotators to ”always consult the AI” or ”edit your code repeatedly,” as this would deviate from real-world scenarios and undermine our intention to use human-annotated data. Considering these reasons, we did not collect programming traces for the entire test set. While we still hope that the number of samples of four different combinations is at least balanced. At this stage, the number of samples for combinations involving all four data types was relatively similar. So we asked annotators to label additional programming process traces for combinations with fewer samples and collected the corresponding traces. Meanwhile, for combinations with slightly more samples, we discarded some of their traces. Through this process, we established our final benchmark. Simplified examples of the annotated data is illustrated in Figure 13.
Figure 13: Simplified examples of APEval, which covering various code editing scenarios that require integrating multiple types of information to infer user intent. The left example checks if any two numbers in a list are closer than a given threshold. The current logic is flawed and should verify if the absolute difference between two values is less than $t$. The model must detect this issue, fix the error, and generate the remaining code. The right example shows a programmer replacing incorrect code with a corrected version. Without historical edits, the model cannot infer the function’s intent. Thus, it must use edit history to make accurate code edits.
Appendix D Additional details about Programming-Instruct
@@
In addition to generating H and U, as discussed in Section 4.2, we also simulate the programmer’s specification of the target area and model interactions in a chat-style format. The target modification area is created using a random algorithm, as described in Appendix E, while the chat-style interaction is generated using LLMs which is similar to the generation of instructions. Prompts used for it are provided in Appendix L.
Appendix E Target area representation
@@
Appendix F Discussion about thought process
@@
Appendix G Conversation retrieval for Assistant-Conversation
@@
Appendix H Evaluation results of other benchmarks
@@
Appendix I Additional evaluation results on APEval
@@
Appendix J Chat template
@@
Appendix K Prompts for evaluation
@@
Appendix L Prompts for data collection
We design specific system prompts and few-shot examples to collect high-quality training data, as we find that many examples are very difficult to complete with current LLMs, and only a few of them can be successfully completed using rough prompts. For AIprogrammer, we utilize LLMs to simulate programmers at three different skill levels, with each level using a distinct set of prompts as shown in Tables 7, 8 and 9. Additionally, prompts used for evaluating whether the outputs align with user intent, generating user instructions, and facilitating chat interactions between models and users are outlined in Tables 12, 10 and 11. Partial few-shot examples are shown in Figures 19, 20, 21, 22, 23 and 24.
Appendix M Limitations and future work